OpenStudy (anonymous):

find the following indefinite integral using u-substitution to solve... HELP

OpenStudy (anonymous):

\[\int\limits_{}^{}6x \sqrt{x^2-25}dx\]

OpenStudy (anonymous):

I know I let u=x^2-25 which makes du=2xdx

OpenStudy (anonymous):

i am stuck at a particular step.. i will show you what i have so far

OpenStudy (mathmate):

hint: try u=sqrt(x^2-25) or try differentiating sqrt(x^2-25) and see what you get.

OpenStudy (dumbcow):

ok then dx = du/2x \[\int\limits_{}^{}\frac{6x \sqrt{u}}{2x} du =3 \int\limits_{}^{}\sqrt{u} du\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\sqrt{x^2-25}\times6xdx = 3\int\limits_{}^{}\sqrt{x^2-25}\times2xdx = 3\int\limits_{}^{}\sqrt{u}du\]

OpenStudy (anonymous):

i'm lost after that

OpenStudy (dumbcow):

\[\sqrt{u} = u^{1/2}\] use the power rule

OpenStudy (dumbcow):

\[\large \int\limits_{}^{}x^{n} = \frac{x^{n+1}}{n+1}\]

OpenStudy (anonymous):

yeah so i would fill fill in x^2 -25 now for that step or no?

OpenStudy (dumbcow):

no leave it in terms of u until you have finished integrating, then the last step will be to substitute the x^2 -25 back in for u

OpenStudy (anonymous):

ok so i got 2u^(3/2)+c ... is that correct?

OpenStudy (dumbcow):

yep

OpenStudy (anonymous):

ok so then substitute in now? giving me.. 2(x^2-25)^(3/2) + c ..?

OpenStudy (dumbcow):

yes you could also write it as ...2sqrt(x^2-25)(x^2 -25) + c

OpenStudy (anonymous):

ok so then my final answer is \[2x^2-50\sqrt{x^2-25}+c\]..? or just leave it \[2(x^2-25)\sqrt{x^2-25}+c\]?

OpenStudy (dumbcow):

i would just leave it in factored form, but both are correct

OpenStudy (anonymous):

thank you so much!